# Federated Large Language Models: Current Progress and Future Directions

### v1

@@ 2.5 Evaluation
Datasets FederatedScope-LLM [31] is a comprehensive fine-tuning package for diver federated settings, including various training and evaluation datasets. In these settings, communication and computation costs, as well as memory and multi-GPU efficiency are considered. OpenFedLLM [85] is another federated LLM training framework, where clients are trained on underutilized distributed private data and can collaboratively train a shared model without data transmission.
Benchmarks FedLLM-Bench [84] focuses on providing real user-annotated datasets in several settings including, multilingual collaborative learning, multi-user preferences, and conversations from different chat bots. Fedmlsecurity [26] on the other hand focuses more on the attacks and defenses scenarios in federated learning. This benchmark provides various backbone model implementations and different attacks and defense configurations. Profit [14] benchmarks the trade-off between personalization and robustness in federated learning. The benchmarking results show robust performance when using a small learning rate with many local epochs for personalization, especially with an adaptive client-side optimizer.
Convergence Analysis FedPEAT [13] proposes a server-device collaborative learning framework, where parameter-efficient learning methods and an emulator are incorporated. The convergence analysis in FedPEAT is comprehensive to demonstrate that it’s 4.6 times faster than the conventional Fed-FT and uses 4.9 times less communications compared with it. FedMeZO [36] integrates a zeroth-order optimization method in federated LLM learning, which converges faster than conventional methods and also significantly reduces GPU memory usage.
@@
Prompt Generation FedTPG [47], TPFL [102]
Few-Shot Scenario FedFSL [3]
Chain of Thoughts Fed-SP-SC [39], FedLogic [78]
Personalization pFL [22], FedLogic [82], Fed-DPT  [67]
Multi-Domain FedAPT [51], Profit [14]
Parameter Efficient FedPepTAO [5], FedLoRA [77]
Communication Efficient FedPrompt [98]
Blackbox FedBPT [53]
Retrieval Augmented Generation FeB4RAG [64]
@@
Multilingual [101], Recommender Systems [21, 99],
Medical VQA [105], Weather Forecasting [11],
Virtual Reality [104]
Table 3: Overview of Topics in Prompt Learning
3 Prompt Learning
Prompt-based learning involves creating a prompting function to achieve optimal performance on a task [37]. This requires either a human or an algorithm to choose the best template for each task. In FL, large LLMs incur high communication costs and need significant local data for training. Prompt learning, which fine-tunes soft prompts without altering LLMs, emerges as a promising method to reduce communication costs, as shown in Table 3.
@@ 3.1 Prompt Generation
PromptFL [23] emphasizes training prompts instead of a shared model, using foundation models like CLIP for efficient global aggregation and local training, even with limited data. This boosts privacy, efficiency, and performance in FL environments.  [47] introduces a context-aware prompt generation network based on task-related textual inputs, improving model adaptability to diverse classes across remote clients.  [102] presents Twin Prompt Federated Learning (TPFL), which integrates visual and textual modalities to overcome the limitations of single-modality prompt tuning in federated learning, improving data representation across nodes.
@@ 3.2 Few-shot Scenario
LLMs are fine-tuned within a federated, privacy-preserving framework for various language tasks, relying on hundreds of thousands of labeled samples from mobile clients. However, mobile users often resist or lack expertise in data labeling, leading to a few-shot scenario[3] introduce FeS, the first framework for Federated Few-Shot Learning in mobile NLP. FeS incorporates curriculum pacing with pseudo labels, representational diversity for sample selection, and co-planning of model layer depth and capacity. This reduces training time, energy, and network traffic, enabling efficient fine-tuning on resource-limited devices.
@@ 3.3 Chain of Thoughts
Recent research shows that Chain of Thought (CoT), a series of reasoning steps, greatly improves LLMs’ complex reasoning abilities [68]. Fed-SP-SC [39] improve LLM accuracy by using federated prompting with CoT reasoning, which employs a majority voting system on synonymous questions to boost zero-shot reasoning. FedLogic [78] enhances LLM interpretability and effectiveness through optimized CoT prompt selection across domains, using variational expectation maximization (V-EM) for cross-domain personalization. These advancements improve LLM adaptability and reasoning capabilities.
@@ 3.4 Personalization
Inter-client uncertainty refers to differences between users’ local models within the same time frame. Higher uncertainty signals more diverse local data distributions among devices, emphasizing the need for personalization. [22] introduce a prompt-based personalized FL (pFL) method for medical VQA, addressing data heterogeneity and privacy by treating distinct organ datasets as separate clients. This method uses pFL to develop unique transformer-based VQA models per client while reducing the communication burden via compact prompt parameters. A reliability parameter prevents performance loss and irrelevant contributions. [82] propose FedLogic, integrating FL with prompt selection optimization for LLMs to enhance chain-of-thought (CoT) reasoning across domains while maintaining privacy. FedLogic balances generality and personalization in CoT prompts by exchanging knowledge between models, fine-tuning rule learning, and optimizing performance through a bi-level program with KL-divergence constraints. [67] propose Fed-DPT, which leverages pre-trained visual language models and dual prompt tuning to improve federated learning in scenarios with domain-diverse client data.
@@ 3.5 Multi-domain
Su et al. [51] introduce FedAPT, which integrates FL with adaptive prompt tuning to tackle multi-domain collaborative learning. FedAPT personalizes prompts for each client’s domain while enabling collaboration across clients without sharing raw data. Experiments show that FedAPT outperforms existing methods in handling domain shifts. [14] address the gap by evaluating FedAvg and FedSGD with personalization in the context of prompt tuning for LLMs. Their benchmark highlights the robustness of federated-trained prompts with small learning rates and many local epochs, showing that regularization and prompt interpolation improve the personalization-robustness trade-off in limited computation settings.
@@ 3.6 Parameter Efficient
Prompt Tuning [5] puts forth a novel Parameter-efficient prompt Tuning approach, FedPepTAO, encompassing Adaptive Optimization to facilitate efficient, effective FL of LLMs. Firstly, a unique and efficient partial prompt tuning method is proposed to simultaneously optimize both performance and efficiency. Secondly, a novel adaptive optimization technique is introduced to combat client drift issues across both the device and server sides and thus enhance overall performance.
LowRank Adaptation Inspired by recent advancements in fine-tuning LLMs, particularly the LowRank Adaptation (LoRA) technique, which preserves general linguistic knowledge in a pre-trained full-rank model while retaining domain-specific knowledge in a low-rank matrix, [77] propose FedLoRA for PFL. FedLoRA aims to maintain shared general knowledge in a common full-rank matrix and capture client-specific knowledge in a personalized low-rank matrix. However, unlike LoRA, FedLoRA’s full-rank matrix requires training from scratch, which is affected by data heterogeneity. To address this, they propose a new training strategy to mitigate these effects.
@@ 3.7 Communication Efficient
Zhao et al. [98] introduce FedPrompt, which combines FL with prompt tuning to tackle challenges in decentralized data processing, prioritizing communication efficiency and privacy. Their method uses a split aggregation framework, significantly reducing communication costs by transmitting only 0.01% of a pre-trained language model’s parameters, with minimal accuracy loss on both IID and Non-IID data. FedPrompt enhances FL efficiency while maintaining strong privacy protections, making large language models more practical for decentralized applications.
@@ 3.8 Blackbox
Applying FL to black-box LLMs faces significant challenges, including restricted access to model parameters, high computational demands, and substantial communication overheads. To address these issues, Sun et al. [53] introduce Federated Black-box Prompt Tuning (FedBPT), an innovative framework that overcomes these obstacles. FedBPT’s key advantage lies in its ability to operate without client access to model parameters. The framework focuses on training optimal prompts and employs gradient-free optimization methods, effectively reducing the number of exchanged variables. This approach significantly enhances communication efficiency and minimizes both computational and storage costs.
@@ 3.9 Applications
Researchers have explored various innovative approaches to applying FL in LLM-based systems across diverse domains, including multilingual processing, recommender systems, medical imaging, weather forecasting, and virtual reality services. These studies demonstrate how FL can be combined with other advanced techniques such as prompt tuning, personalization strategies, and domain-specific adaptations to enhance the performance and applicability of LLMs in real-world scenarios while maintaining data confidentiality and addressing the unique challenges of distributed learning environments.
Multilingual: Zhao et al. [101] address the challenges in fine-tuning multilingual LLMs, particularly for low-resource languages, where data-sharing restrictions and linguistic differences create significant barriers. To overcome these obstacles, it proposes a novel approach called Multilingual Federated Prompt Tuning. This paradigm leverages FL to enable parameter-efficient fine-tuning for multilingual tasks across various regions or devices. The process involves each country fine-tuning the model locally and then sharing the updated parameters with a central server for aggregation. This method results in a global model that capitalizes on collective knowledge and exposes the system to a broader range of linguistic patterns. To account for the diverse linguistic characteristics across countries, the authors implement prompt encoders that enhance the model’s ability to generalize and adapt to different languages. This approach aims to democratize access to the benefits of LLMs for users of various languages, especially those in low-resource regions, while respecting data privacy concerns and accommodating linguistic diversity.
Recommender Systems: In the domain of recommendation systems, Cross-Domain Recommendation (CDR) poses notable challenges concerning data privacy breaches and inefficient transfer of knowledge within existing Factorization Machine-based approaches. To mitigate these concerns, Guo et al. [21] have proposed the Privacy-Preserving Federated Content Representation (PFCR) framework. This framework incorporates a FL architecture with local client interactions and secure gradient encryption mechanisms, thereby addressing privacy issues. Besides, it exploits textual descriptions to project items into a shared feature space and enhances content representations through federated learning strategies, coupled with prompt fine-tuning methods, to facilitate more effective cross-domain knowledge transfer.
Medical VQA: Zhu et al. [105] introduce a novel prompt-based personalized federated learning (pFL) method for medical visual question answering (VQA), addressing data heterogeneity and privacy concerns. Their approach treats datasets from different organs as separate clients, using small learnable prompts for efficient information sharing between clients. The method also incorporates a reliability parameter to mitigate negative effects from low-performing or irrelevant clients, enhancing the overall robustness of the federated learning process.
Weather Forecasting: Chen et al.[11] develops a cross-regional foundation model for weather forecasting using complex meteorological data. To address data privacy concerns, they propose a novel federated learning approach for collaboratively training a spatio-temporal Transformer-based model with heterogeneous data from multiple participants. The study also introduces a prompt learning mechanism to accommodate low-resourced sensors’ constraints, enhancing the model’s adaptability and efficiency in areas with limited technological infrastructure.
Virtual Reality: Zhou et al. [104] tackles the challenge of resource allocation for customized virtual reality (VR) services in mobile edge computing (MEC) systems serving heterogeneous users. They propose the FedPromptDT framework, which combines federated learning (FL) and prompt-based sequence modeling. This approach pre-trains a FedPromptDT model using FL, allowing for collaborative learning across distributed clients while preserving data privacy. The framework’s key innovation is its incorporation of prompt design, enabling the model to effectively process environmental information and user-preferred allocation decisions. This design allows the FedPromptDT model to easily adapt to different user environments without retraining, resulting in efficient and personalized resource allocation decisions. Consequently, the system enhances the overall performance and user experience of customized VR services in MEC systems, addressing the unique needs of heterogeneous users.
4 Potential Directions
@@
Federated Inference Real-time On-device Inference
@@ 4.1 Real World Deployment
One of the critical challenges in deploying federated LLMs to ensure that these models can operate effectively in real-world scenarios, where the data is often heterogeneous and the computing resources are often constrained. Recent works have introduced techniques like LoRA within personalized federated learning to address these challenges by maintaining shared general knowledge while capturing client-specific adaptations, thereby mitigating the impact of data heterogeneity [77].
However, despite these advances, optimizing federated learning frameworks for deployment in more diverse real-world environments remains a challenge. For example, developing personalized AI agents that can fine-tune on confidential data while respecting user privacy will require innovations in privacy-preserving techniques to enable collaborative learning with large scale models. Additionally, while efficient model adaptation will be critical in enhancing the performance and scalability of LLMs in decentralized settings [85], current approaches often make trade-offs between model accuracy and resource efficiency. Furthermore, robust deployment strategies that can handle the variability in real-world data and infrastructure, particularly in resource-contrained environments, may also benefit from techniques such as swarm intelligence [48] that distribute computational tasks across multiple agents.
@@ 4.2 Multimodality Models
Integrating multimodal data into federated LLMs presents another opportunity for enhancing the versatility and applicability of these models. Methods such as FedClip [42] have shown potential in combining visual and textual data to improve performance across various tasks. However, how to ensuring seamless integration of different data modalities while maintaining data privacy and model efficiency poses new challenges. Recent studies such as [80] have noted the complexity of aligning diverse data modalities in decentralized settings, which often results in sub-optimal performance and increased computational costs. Future research may explore the co-optimization of multiple models handling different modalities, focusing on reducing these inefficiencies and developing robust frameworks for more comprehensive and context-aware AI systems in federated environments.
@@ 4.3 Federated Pre-Training
Given the high computational costs associated with pre-training LLMs, there is a growing need for strategies that reduce these costs without sacrificing model performance. Federated pre-training [58, 70] offers a potential solution by enabling models to be adapted to specific domains without relying on extensive centralized datasets. Among the challenges in federated pre-training, the efficiency of data exchange protocols remains a crucial factor. Recent strategies like fully-sharded data parallelism [49], which aims to minimize memory consumption and optimize communication between GPUs during training, have made federated pre-training more viable even with limited hardware resources. Future research may further explore these techniques to strike a better balance between computational costs and the benefits of federated pre-training, possibly by optimizing data exchange protocols or developing more efficient model architectures.
4.4 Federated Inference on Local Data

### v2

@@ 2.5 Evaluation
Datasets FederatedScope-LLM [kuang2023federatedscope] is a comprehensive fine-tuning package for diverse federated settings, including various training and evaluation datasets. In these settings, communication and computation costs, as well as memory and multi-GPU efficiency, are considered. OpenFedLLM [ye2024openfedllm] is another federated LLM training framework, where clients are trained on underutilized distributed private data and can collaboratively train a shared model without data transmission.
Benchmarks FedLLM-Bench [ye2024fedllm] focuses on providing real user-annotated datasets in several settings, including multilingual collaborative learning, multi-user preferences, and conversations from different chatbots. Fedmlsecurity [han2023fedmlsecurity], on the other hand, focuses more on the attacks and defense scenarios in federated learning. This benchmark provides various backbone model implementations and different attacks and defense configurations. Profit [collins2023profit] benchmarks the trade-off between personalization and robustness in federated learning. The benchmarking results show robust performance when using a small learning rate with many local epochs for personalization, especially with an adaptive client-side optimizer.
Convergence Analysis FedPEAT [chua2023fedpeat] proposes a server-device collaborative learning framework, where parameter-efficient learning methods and an emulator are incorporated. The convergence analysis in FedPEAT is comprehensive to demonstrate that it’s 4.6 times faster than the conventional Fed-FT and uses 4.9 times less communications compared with it. FedMeZO [ling2024convergence] integrates a zeroth-order optimization method in federated LLM learning, which converges faster than conventional methods and also significantly reduces GPU memory usage.
3 Prompt Learning
Prompt-based learning involves creating a prompting function to achieve optimal performance on a task [liu23survey]. This requires either a human or an algorithm to choose the best template for each task. In FL, large LLMs incur high communication costs and need significant local data for training. Prompt learning, which fine-tunes soft prompts without altering LLMs, emerges as a promising method to reduce communication costs, as shown in Table 3.
Table 3: Overview of Topics in Prompt Learning.
@@ 3 Prompt Learning
Prompt Generation FedTPG [qiu2023text], TPFL [zhao2023inclusive]
Few-Shot Scenario FedFSL [cai2023federated]
Chain of Thoughts Fed-SP-SC [liu2023federated], FedLogic [xing2023fedlogic]
Personalization pFL [guo2023pfedprompt], FedLogic [yang2023efficient], Fed-DPT  [wei2023dual]
Multi-Domain FedAPT [su2022federated], Profit [collins2023profit]
Parameter Efficient FedPepTAO [che2023federated], FedLoRA [wu2023fedlora]
Communication Efficient FedPrompt [zhao2023fedprompt]
Blackbox FedBPT [sun2023fedbpt]
Retrieval Augmented Generation FeB4RAG [wang2024feb4rag]
@@
Multilingual [zhao2023breaking], Recommender Systems [guo2024prompt, zhao2024llm],
Medical VQA [zhu2024prompt], Weather Forecasting [chen2023prompt],
Virtual Reality [zhou2024federated]
@@ 3.1 Prompt Generation
PromptFL [guo2023promptfl] emphasizes training prompts instead of a shared model, using foundation models like CLIP for efficient global aggregation and local training, even with limited data. This boosts privacy, efficiency, and performance in FL environments.  [qiu2023text] introduces a context-aware prompt generation network based on task-related textual inputs, improving model adaptability to diverse classes across remote clients.  [zhao2023inclusive] presents Twin Prompt Federated Learning (TPFL), which integrates visual and textual modalities to overcome the limitations of single-modality prompt tuning in federated learning, improving data representation across nodes.
@@ 3.2 Few-shot Scenario
LLMs are fine-tuned within a federated, privacy-preserving framework for various language tasks, relying on hundreds of thousands of labeled samples from mobile clients. However, mobile users often resist or lack expertise in data labeling, leading to a few-shot scenario[cai2023federated] introduce FeS, the first framework for Federated Few-Shot Learning in mobile NLP. FeS incorporates curriculum pacing with pseudo labels, representational diversity for sample selection, and co-planning of model layer depth and capacity. This reduces training time, energy, and network traffic, enabling efficient fine-tuning on resource-limited devices.
@@ 3.3 Chain of Thoughts
Recent research shows that Chain of Thought (CoT), a series of reasoning steps, greatly improves LLMs’ complex reasoning abilities [jason2022cot]. Fed-SP-SC [liu2023federated] improve LLM accuracy by using federated prompting with CoT reasoning, which employs a majority voting system on synonymous questions to boost zero-shot reasoning. FedLogic [xing2023fedlogic] enhances LLM interpretability and effectiveness through optimized CoT prompt selection across domains, using variational expectation maximization (V-EM) for cross-domain personalization. These advancements improve LLM adaptability and reasoning capabilities.
@@ 3.4 Personalization
Inter-client uncertainty refers to differences between users’ local models within the same time frame. Higher uncertainty signals more diverse local data distributions among devices, emphasizing the need for personalization. [guo2023pfedprompt] introduce a prompt-based personalized FL (pFL) method for medical VQA, addressing data heterogeneity and privacy by treating distinct organ datasets as separate clients. This method uses pFL to develop unique transformer-based VQA models per client while reducing the communication burden via compact prompt parameters. A reliability parameter prevents performance loss and irrelevant contributions. [yang2023efficient] propose FedLogic, integrating FL with prompt selection optimization for LLMs to enhance chain-of-thought (CoT) reasoning across domains while maintaining privacy. FedLogic balances generality and personalization in CoT prompts by exchanging knowledge between models, fine-tuning rule learning, and optimizing performance through a bi-level program with KL-divergence constraints. [wei2023dual] propose Fed-DPT, which leverages pre-trained visual language models and dual prompt tuning to improve federated learning in scenarios with domain-diverse client data.
@@ 3.5 Multi-domain
Su et al. [su2022federated] introduce FedAPT, which integrates FL with adaptive prompt tuning to tackle multi-domain collaborative learning. FedAPT personalizes prompts for each client’s domain while enabling collaboration across clients without sharing raw data. Experiments show that FedAPT outperforms existing methods in handling domain shifts. [collins2023profit] address the gap by evaluating FedAvg and FedSGD with personalization in the context of prompt tuning for LLMs. Their benchmark highlights the robustness of federated-trained prompts with small learning rates and many local epochs, showing that regularization and prompt interpolation improve the personalization-robustness trade-off in limited computation settings.
@@ 3.6 Parameter Efficient
Prompt Tuning [che2023federated] puts forth a novel Parameter-efficient prompt Tuning approach, FedPepTAO, encompassing Adaptive Optimization to facilitate efficient, effective FL of LLMs. Firstly, a unique and efficient partial prompt tuning method is proposed to simultaneously optimize both performance and efficiency. Secondly, a novel adaptive optimization technique is introduced to combat client drift issues across both the device and server sides and thus enhance overall performance.
LowRank Adaptation Inspired by recent advancements in fine-tuning LLMs, particularly the LowRank Adaptation (LoRA) technique, which preserves general linguistic knowledge in a pre-trained full-rank model while retaining domain-specific knowledge in a low-rank matrix, [wu2023fedlora] propose FedLoRA for PFL. FedLoRA aims to maintain shared general knowledge in a common full-rank matrix and capture client-specific knowledge in a personalized low-rank matrix. However, unlike LoRA, FedLoRA’s full-rank matrix requires training from scratch, which is affected by data heterogeneity. To address this, they propose a new training strategy to mitigate these effects.
@@ 3.7 Communication Efficient
Zhao et al. [zhao2023fedprompt] introduce FedPrompt, which combines FL with prompt tuning to tackle challenges in decentralized data processing, prioritizing communication efficiency and privacy. Their method uses a split aggregation framework, significantly reducing communication costs by transmitting only 0.01% of a pre-trained language model’s parameters, with minimal accuracy loss on both IID and Non-IID data. FedPrompt enhances FL efficiency while maintaining strong privacy protections, making large language models more practical for decentralized applications.
@@ 3.8 Blackbox
Applying FL to black-box LLMs faces significant challenges, including restricted access to model parameters, high computational demands, and substantial communication overheads. To address these issues, Sun et al. [sun2023fedbpt] introduce Federated Black-box Prompt Tuning (FedBPT), an innovative framework that overcomes these obstacles. FedBPT’s key advantage lies in its ability to operate without client access to model parameters. The framework focuses on training optimal prompts and employs gradient-free optimization methods, effectively reducing the number of exchanged variables. This approach significantly enhances communication efficiency and minimizes both computational and storage costs.
@@ 3.9 Applications
Researchers have explored diverse applications of FL in LLM-based systems across domains such as multilingual processing, recommender systems, medicine, weather forecasting, and virtual reality. These studies illustrate how FL can be integrated with techniques including prompt tuning, personalization, and domain-specific adaptations to improve model performance while preserving data privacy and addressing challenges inherent to distributed environments.
Multilingual: Zhao et al. [zhao2023breaking] propose Multilingual Federated Prompt Tuning to enable parameter-efficient adaptation of multilingual LLMs without sharing raw data. Local prompt encoders capture region-specific linguistic characteristics, and aggregated updates produce a global model that better generalizes to low-resource languages while respecting data restrictions.
Recommender Systems: Guo et al. [guo2024prompt] introduce the Privacy-Preserving Federated Content Representation (PFCR) framework, which combines FL with secure gradient encryption and prompt-based content representation. By leveraging textual item descriptions and shared feature spaces, PFCR enhances cross-domain recommendation while mitigating privacy risks.
Medical VQA: Zhu et al. [zhu2024prompt] present a personalized prompt-based FL approach for medical VQA, treating datasets from different organs as separate clients. Lightweight prompts enable efficient knowledge sharing, while a reliability weighting mechanism reduces negative influence from low-quality clients.
Weather Forecasting: Chen et al. [chen2023prompt] develop a federated spatio-temporal foundation model for cross-regional weather forecasting. Prompt learning improves adaptation to heterogeneous and low-resource sensors, enabling collaborative forecasting without compromising sensitive meteorological data.
Virtual Reality: Zhou et al. [zhou2024federated] propose FedPromptDT, integrating FL and prompt-based sequence modeling for personalized resource allocation in mobile edge VR services. Prompt design allows rapid adaptation to varying environments and user preferences without retraining, improving efficiency and user experience.
4 Emerging Directions
@@
Federated AI Agents Federated Inference on Local Data
Federated Agents across Clients
@@ 4.1 Real World Deployment
One of the critical challenges in deploying federated LLMs to ensuring that these models can operate effectively in real-world scenarios, where the data is often heterogeneous and the computing resources are often constrained. Recent works have introduced techniques like LoRA within personalized federated learning to address these challenges by maintaining shared general knowledge while capturing client-specific adaptations, thereby mitigating the impact of data heterogeneity [wu2023fedlora].
However, despite these advances, optimizing federated learning frameworks for deployment in more diverse real-world environments remains a challenge. For example, developing personalized AI agents that can fine-tune on confidential data while respecting user privacy will require innovations in privacy-preserving techniques to enable collaborative learning with large-scale models. Additionally, while efficient model adaptation will be critical in enhancing the performance and scalability of LLMs in decentralized settings [ye2024openfedllm], current approaches often make trade-offs between model accuracy and resource efficiency. Furthermore, robust deployment strategies that can handle the variability in real-world data and infrastructure, particularly in resource-constrained environments, may also benefit from techniques such as swarm intelligence [qu2024federated] that distribute computational tasks across multiple agents.
@@ 4.2 Multimodality Models
Integrating multimodal data into federated LLMs presents another opportunity for enhancing the versatility and applicability of these models. Methods such as FedClip [lu2023fedclip] have shown potential in combining visual and textual data to improve performance across various tasks. However, how to ensuring seamless integration of different data modalities while maintaining data privacy and model efficiency poses new challenges. Recent studies, such as [xu2024survey] have noted the complexity of aligning diverse data modalities in decentralized settings, which often results in sub-optimal performance and increased computational costs. Future research may explore the co-optimization of multiple models handling different modalities, focusing on reducing these inefficiencies and developing robust frameworks for more comprehensive and context-aware AI systems in federated environments.
@@ 4.3 Federated Pre-Training
Given the high computational costs associated with pre-training LLMs, there is a growing need for strategies that reduce these costs without sacrificing model performance. Federated pre-training [tian2022fedbert, wen2024pre] offers a potential solution by enabling models to be adapted to specific domains without relying on extensive centralized datasets. Among the challenges in federated pre-training, the efficiency of data exchange protocols remains a crucial factor. Recent strategies like fully-sharded data parallelism [sani2024future], which aims to minimize memory consumption and optimize communication between GPUs during training, have made federated pre-training more viable even with limited hardware resources. Future research may further explore these techniques to strike a better balance between computational costs and the benefits of federated pre-training, possibly by optimizing data exchange protocols or developing more efficient model architectures.
4.4 Federated AI Agents